Solve the equation check for extraneous solutions
9|9-8x|=2x+3

Found 2 solutions by josgarithmetic, It is costly:
Answer by josgarithmetic(39617)

Either 9(9-8x)=2x+3
81-72x=2x+3
81=74x+3
74x=78
x=78/74
x=39/37

OR

9(8x-9)=2x+3
72x-81=2x+3
70x=3+81
70x=84
x=84/70
x=12/10
x=6/5

Answer by It is costly(175)